OpenStudy (anonymous):

2/a^2+4a+4 + 9/a^2-4

OpenStudy (anonymous):

Do you mean 2/(a^2+4a+4) + 9/(a^2-4)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The two denominators can be factored. Give it a try.

OpenStudy (anonymous):

2/(a+4) (a+1) + 9/(a+4)(a-4)

OpenStudy (anonymous):

Close. Did you actually multiply them out to see if you get back the original expression?

OpenStudy (anonymous):

u r doin silly mistake

OpenStudy (anonymous):

i figured out what i did wrong but i am having trouble finding the lcd

OpenStudy (anonymous):

u can... . try...

OpenStudy (anonymous):

ur tryin has much more importance

OpenStudy (anonymous):

i think its (a+2)(a+2) (a+4)(a-4)

OpenStudy (anonymous):

no....

OpenStudy (anonymous):

second part is wrong

OpenStudy (anonymous):

can u just tell me how to do it then

OpenStudy (anonymous):

First correctly factor the denominators in 2/(a^2+4a+4) + 9/(a^2-4)

OpenStudy (anonymous):

Ooops, missed an earlier post of yours.

OpenStudy (anonymous):

(a+4)(a-4) is not correct.

OpenStudy (anonymous):

ya.. just examine that part

OpenStudy (anonymous):

i no i fixed it

OpenStudy (anonymous):

(a+2) (a+2)

OpenStudy (anonymous):

u can write (a^2-4) as (a^2-2^2) =(a-2)(a+2)

OpenStudy (anonymous):

got it ?????

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

is this right so far 2a-4/(a+2) (a+2) (a-2) + 9a+18/(a+2 (a+2) (a-2)

OpenStudy (anonymous):

rite

OpenStudy (anonymous):

2/(a-2)(a-2) + 9/(a+2)(a-2)

OpenStudy (anonymous):

a^2+4a+4 =(a+2)(a+2)

OpenStudy (anonymous):

its +4 my friend

OpenStudy (anonymous):

correction : 2/(a+2)(a_2) + 9/(a+2)(a-2)

OpenStudy (anonymous):

Dumb computer : 2/(a+2)(a+2) + 9/(a+2)(a-2)

OpenStudy (anonymous):

now its rite

OpenStudy (anonymous):

Yes, finally. lol

OpenStudy (anonymous):

final answer 7a+14/(a+2)a+2)(a-2)

OpenStudy (anonymous):

2a+9a=11a

OpenStudy (anonymous):

i got anther problem 3/a^2+12a+36 + 5/a^2-36

OpenStudy (anonymous):

same type as before

OpenStudy (anonymous):

No, final answer is not correct.

OpenStudy (anonymous):

3/(a+6)(a+6) + 5 /(a+6)(a-6)

OpenStudy (anonymous):

Working with this latest problem, what is the LCD?

OpenStudy (anonymous):

with the problem im doing now

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

i have (a+6)(a+6)(a-6)

OpenStudy (anonymous):

Yes. So now change the two fractions so they both have this same denominator.

OpenStudy (anonymous):

so i have 3a-18/(a+6)(a+6)(a-6) + 5a+30/(a+6)(a+6)(a-6)

OpenStudy (anonymous):

That looks correct. Now combine the numerators.

OpenStudy (anonymous):

8a+12

OpenStudy (anonymous):

Yes. Of course, that goes over the denominator. 8a +12 does not factor in any way that can cancel with something in the denominator, so that should be the final answer.

OpenStudy (anonymous):

(8a +12)/(a+6)(a+6)(a-6)